The pdfpages Package^{∗}

Andreas MATTHIAS

                      amat@kabsi.at

2002/12/14

Abstract

This package makes it easy to insert pages of external PDF documents. It is based on pdfL^{A}TEX and does not work with L^{A}TEX.

Contents

1 Introduction 1

2 Usage 2 2.1 Package Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 The Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.4 Pitfalls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3 Required Packages 9

4 Acknowledgment 10

1 Introduction

Creating PDF documents, it is sometimes useful to insert pages of other, external PDF documents. This can be done with the \includegraphics command from

the graphics package. But a simple \includegraphics{doc.pdf} normally pro-

duces ‘Overfull \hbox’ and ‘Overfull \vbox’ warnings, because the size of the inserted pages does not match the print space.

The pdfpages package makes it easy to insert pages of external PDF documents without worrying about the print space. Here are some features of the pdfpages package: Several logical pages can be arranged onto each sheet of paper and the layout can be changed individually. A lot of hypertext operations are supported, like links to the inserted pages, links to the original PDF document, threads, etc.

∗This file has version number v0.2l, last revised 2002/12/14.

2 Usage

2.1 Package Options

\usepackage[⟨options⟩]{pdfpages}

⟨option⟩ – final: Inserts pages. This is the default. draft: Does not insert pages, but prints a box and the filename instead. enable-survey: Activates survey functionalities. (ex- perimental, subject to change)

2.2 Commands

\includepdf Inserts pages of an external PDF document.

\includepdf[⟨key=val⟩]{⟨filename⟩}

⟨key=val⟩ – A comma separated list of options using the

⟨key⟩=⟨value⟩ syntax.

⟨filename⟩ – Filename of the PDF document. (The filename must not contain any blanks!)

The following list describes all possible options of \includepdf. All options are using the ⟨key=value⟩ syntax.

• Main options:

pages Selects pages to insert. The argument is a comma separated list, containing page numbers (pages={3,5,6,8}), ranges of page numbers (pages={4-9}) or any combination. To insert empty pages use {}. E.g.: pages={3,{},8-11,15} will insert page 3, an empty page, and pages 8, 9, 10, 11, and 15. Page ranges are specified by the following syntax: ⟨m⟩-⟨n⟩. This selects all pages between ⟨m⟩ and ⟨n⟩. Omitting ⟨m⟩ defaults to the first page, omitting ⟨n⟩ defaults to the last page of the document. E.g.: pages=- will insert all pages of the document. (Default: pages=1)

nup Puts multiple logical pages onto each sheet of paper. The syntax of this option is: nup=⟨xnup⟩x⟨ynup⟩. Where ⟨xnup⟩ and ⟨ynup⟩ specify the number of logical pages in horizontal and vertical direction, which are arranged on each sheet of paper. (Default: nup=1x1)

landscape Specifies the format of the sheet of paper, which is rotated by 90 degrees. This does not affect the logical pages, which will not be rotated by the ‘landscape’ option. To rotate the logical pages use the ‘angle’ option (e.g. ‘angle=90’). Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: landscape=false)

• Layout options:

delta Puts some horizontal and vertical space between the logical pages. The argument should be two dimensions, separated by space. See Chapter 2.3 and Figure 1. (Default: delta=0 0).

offset Displaces the origin of the inserted pages. The argument should be two dimensions, separated by space. In ‘oneside’ documents positive values shift the pages to the right and to the top margin, respectively, whereas in ‘twoside’ documents positive values shift the pages to the outer and to the top margin, respectively. See Chapter 2.3 and Figure 1. (Default:

offset=0 0)

frame Puts a frame around each logical page. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: frame=false)

column Pdfpages normally uses ‘row-major’ layout, where successive pages are placed in rows along the paper. The column option changes the output into a ‘column-major’ layout, where successive pages are arranged in columns down the paper. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: column=false)

columnstrict By default the last page is not set in a strict ‘column-major’ layout, if the logical pages do not fill up the whole page. The columnstrict option forces a strict ‘column-major’ layout for the last page. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

columnstrict=false)

1 4 2 5 3

     columnstrict=true

1 3 5 2 4

                             columnstrict=false

openright This option puts an empty page before the first logical page. In combination with nup=2x1, nup=2x2, etc., this means that the first page is on the right side. The same effect can be achieved with the pages option, if an empty page is inserted in front of the first page. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: openright=false)

pagecommand Declares L^{A}TEX commands, which are executed on each sheet

of paper. (Default: pagecommand={\thispagestyle{empty}})

turn By default pages in landscape format are displayed in landscape orientation (if the PDF viewer supports this). With turn=false this can be prohibited. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: turn=true)

noautoscale By default pages are scaled automatically. This can be suppressed with the noautoscale option. In combination with the scale option (from graphicx) you has full control over the scaling process. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

noautoscale=false)

fitpaper Adjusts the paper size to the one of the inserted document. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

fitpaper=false)

doublepages Inserts every page twice. This is useful for 2-up printing, if one wants to cut the stack of paper afterwards to get two copies. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

doublepages=false)

signature Creates booklets by rearranging pages into signatures and setting nup=1x2 or nup=2x1, respectively. This option takes one argument specifying the size of the signature, which should be a multiply of 4. An example for documents in portrait orientation:

\includepdf[pages=-, signature=8,
            landscape]{portrait-doc.pdf}

An example for documents in landscape orientation:

\includepdf[pages=-, signature=8]{landscape-doc.pdf}

• Hypertext options:

link Inserted pages become a target of a hyperlink. The name of the link is ‘⟨filename ⟩.⟨page number ⟩’. The file extension of ⟨filename⟩ must not be stripped. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: link=false)

linkname Changes the default linkname created by the option link. Instead of ⟨filename⟩ the value of this option is used. E.g. linkname=mylink produces the linknames ‘mylink.⟨page number ⟩’.

thread Combines inserted pages to an article thread. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: thread=false)

threadname Several threads are distinguished by their threadnames. By default the threadname is equal to the filename, but it can be changed with this option. This is useful if the same file is inserted twice or more times and should not be combined to one single thread. Or the other way round if pages from different documents should be combined to one single

thread. (Default: threadname=⟨filename ⟩)

linktodoc Lets the inserted pages be hyperlinks to the document from which they were extracted. Note that the PDF-Viewer will not find the file, if ⟨filename⟩ has not filename extension (.pdf). Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default: linktodoc=false)

• Additional hypertext options:

linkfit Specifies the way the viewer displays a linked page. This option changes the default behavior of the option link. Possible values are:

fitb, fith, fitv, fitr, xyz zoom ⟨integer ⟩, etc. These are destina-

tions (\pdfdest) like they are described in [1]. (Default: linkfit=fitr)

linktodocfit By default the option linktodoc opens the page in ‘Fit in Window’ view. Another view can be specified with this option. Possible values are the legal PDF tokens: /FitH ⟨top⟩, /FitV ⟨left⟩, etc. (See [2] for more details.) (Default: linktodocfit=/Fit)

linkfilename Sets the name (with path) of the file to be linked to by the option linktodoc. You will hardly ever need this option. (Default:

linkfilename=⟨filename ⟩)

• Experimental options: (Syntax may change in future versions!)

addtotoc Adds an entry to the table of contents. This option requires five arguments, separated by commas:

addtotoc={⟨page number ⟩,⟨section ⟩,⟨level ⟩,⟨heading ⟩,⟨label ⟩}

⟨page number⟩: Page number of the inserted page. ⟨section⟩: L^{A}TEX sectioning name – e.g., section, subsection, … ⟨level⟩: Number, denoting depth of section – e.g., section=1, subsec-

tion=2, … ⟨heading⟩: Title inserted in the table of contents. ⟨label⟩: Name of the label. This label can be referred to with \ref and

    \pageref.

Note: The order of the five arguments must not be mixed. Otherwise you will get very strange error messages. The addtotoc option accepts multiple sets of the above mentioned five arguments, all separated by commas. The sets must be sorted such that the ⟨page number⟩s are in ascending order. (Strictly speaking they must have the same order as the page numbers specified by the pages option.) The proper recursive definition of the addtotoc option is: addtotoc={⟨toc-list⟩} ⟨toc-list⟩ → ⟨page number⟩,⟨section⟩,⟨level⟩,⟨heading⟩,⟨label⟩[,⟨toc-list⟩]

addtolist Adds an entry to the list of figures, the list of tables, or any other list (e.g. from float.sty). This option requires four arguments, separated by commas:

addtolist={⟨page number ⟩,⟨type ⟩,⟨heading ⟩,⟨label ⟩}

⟨page number⟩: Page number of the inserted page. ⟨type⟩: Name of a floating environment. (figure, table, etc.) ⟨heading⟩: Title inserted into LoF, LoT, etc. ⟨label⟩: Name of the label. This label can be referred to with \ref and

    \pageref.

Like addtotoc, addtolist accepts multiple sets of the above mentioned four arguments, all separated by commas. The proper recursive definition is: addtolist={⟨lof-list⟩} ⟨lof-list⟩ → ⟨page number⟩,⟨type⟩,⟨heading⟩,⟨label⟩[,⟨lof-list⟩]

survey Creates a survey of those pages of the document, which are marked with \AddToSurvey. (\AddToSurvey is a simple command with no arguments. It just writes out labels to the .aux file.) This option may be used when preparing slides to create a survey of only ‘finished’ pages – if pages are build up incrementally. To use this option a special sequence of production steps must be obeyed. Here is a small example:

  --- slides.tex ---
  \documentclass{article}
  \usepackage[draft,enable-survey]{pdfpages}

  \begin{document}
    ... some text ...
    \AddToSurvey
    ... some text ...
    \includepdf[survey,nup=2x2]{slides-tmp.pdf}
  \end{document}

This is the outline of a document, called slides.tex. Run it through pdfL^{A}TEX several times until all cross-references are solved. (L^{A}TEX will produce a warning, if cross-references are not solved, yet.) Now copy the

file slides.pdf to slides-tmp.pdf and rename draft (package option

of pdfpages) to final. The next and final run through pdfL^{A}TEX will actually insert the desired pages, whereas the former runs with draft did just insert blank pages. The inserted pages are hyperlink to the orignial pages. Experienced users would certainly call

  pdflatex '\PassOptionsToPackage{final}{pdfpages} \input{slides}'

instead of exchanging draft and final. Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

survey=false)

survey-nolink Same as option survey except that the inserted pages do not become hyperlinks. This option may be used to create an external survey. To continue the example above (slides.tex), it is now possible to create handouts of the ‘finished’ slides as an external document.

  \documentclass{article}
  \usepackage{xr}
  \externaldocument{slides}
  \usepackage[enable-survey]{pdfpages}
  \begin{document}
    \includepdf[survey-nolink]{slides.pdf}
  \end{document}

Either ‘true’ or ‘false’ (or no value, which is equivalent to ‘true’). (Default:

survey-nolink=false)

xr-prefix Adjusts a prefix to the labels \includepdf is is looking for. The name of the prefix must be the same as the name of the optional argument

of \externaldocument. (Default: xr-prefix={})

Internally the command \includepdf makes use of the \includegraphics command from the graphicx (actually graphics) package. Hence it is possible to use all the options of \includegraphics, too. Options which are not interpreted by

\includepdf are passed directly to \includegraphics.

Especially the ‘trim’ and ‘clip’ options of \includegraphics are quite useful, if only parts of a page should be inserted. (Maybe to cut off the header and footer of the inserted pages.) Just use the ‘trim’ and ‘clip’ options as if they were options of \inlcudepdf. They will be passed to \includegraphics internally.

[figure: delx; delx; dely; dely; offx; offy]

physical sheet of paper  delta=⟨delx⟩ ⟨dely⟩
logical pages  offset=⟨offx⟩ ⟨offy⟩

centered position

Figure 1: Layout \includepdfmerge Inserts pages of several external PDF documents.

\includepdfmerge[⟨key=val⟩]{⟨file-page-list⟩}

⟨key=val⟩ – A comma separated list of options using the

⟨key⟩=⟨value⟩ syntax.

⟨file-page-list⟩ – ⟨filename⟩[,⟨page spec⟩][,⟨file-page-list⟩] A comma separated list of filenames and optional ⟨page spec⟩ specifiers. A ⟨page spec⟩ can be everything the option pages accepts. Leading and trailing spaces of items in the list is stripped.

The \includepdfmerge command uses the same options as \includepdf with one exception. The option pages has no meaning for \includepdfmerge. Instead the ⟨page spec⟩ specifier is used to specify which pages should be inserted. The ⟨page spec⟩ specifier accepts the same values as the pages option. If no ⟨page spec⟩ specifier is given, only the first page will be inserted.

Examples: To create a kind of summary of three PDF documents, it might be nice to insert just the first page of each document and to provide links to the original documents:

   \includepdfmerge[nup=1x3, landscape, linktodoc]
                   {doc1.pdf, doc2.pdf, doc3.pdf}

But sometimes the title page of a document is not the first page. So it would be more pleasant to insert the title page of each document than the first page. This can be done with the ⟨page spec⟩ specifier. The following example inserts the second page of doc1.pdf and the third page of doc2.pdf and doc3.pdf :

   \includepdfmerge[nup=1x3, landscape, linktodoc]
                   {doc1.pdf, 2, doc2.pdf, 3, doc3.pdf, 3}

Here is an example of more complex ⟨page spec⟩ specifiers:

   \includepdfmerge[nup=1x3, landscape, linktodoc]
                   {doc1.pdf, 1-3,
                    doc2.pdf, 3, 5, 9,
                    doc3.pdf, 3-5, 7}

\includepdfset If you need the same options for \includepdf all the time, it is possible to define global options with \includepdfset. The argument of \includepdfset is a comma separated list of options, using the ⟨key⟩=⟨value⟩ syntax. These options are processed each time \includepdf is called. Local options (passed as an optional argument directly to \includepdf) are overwriting global options:

   \includepdfset{⟨global options⟩}
   \includepdf[⟨local options⟩]{pdf-file}

Only options specific to this package can be made global by \includepdfset. Options of the graphicx package are not concerned.

\threadinfodict When using the option thread to create an article thread, it may be useful to create a thread information dictionary, too, which contains informations about the thread, such as its title, author, and creation date. The macro \threadinfodict is used to set these informations. It can be redefined and must contain the thread information dictionary in low-level PDF commands. (See [2] for more information.)

   \renewcommand*{\threadinfodict}
     {/I << /Title (My first thread) /Author (That's me!) >>}

2.3 The Layout

The default layout can be changed by the options delta and offset. Figure 1 shows the meaning of these options.

The inserted logical pages are being centered on the sheet of paper by default. To displace them use the offset option, which argument should be two dimensions. E.g. offset=10mm 14mm means that the logical pages are displaced by 10 mm in horizontal direction and by 14 mm in vertical direction. In ‘oneside’ documents positive values shift the pages to the right and to the top margin, respectively, whereas in ‘twoside’ documents positive values shift the pages to the outer and to the top margin, respectively.

By default logical pages are being arranged side by side. To put some space between them, use the delta option, whose argument should be two dimensions. Figure 1 shows the meaning of delta.

The layout options delta and offset always refer to a sheet of paper in portrait orientation. No matter whether you have set the landscape option to true, or not.

If you are confused about horizontal (x) and vertical (y) directions, just set the option turn=false. Now your PDF viewer shows the pages in the same orientation as in Figure 1. And the options delta and offset have the same meaning as in Figure 1. Regardless of any other options.

2.4 Pitfalls

pagecolor When setting the background color with \pagecolor (a command from color.sty), the first \pagecolor must precede \usepackage{pdfpages}.

       \usepackage{color}
       \pagecolor{white}
       \usepackage{pdfpages}

The color is nonrelevant, it can be changed afterwards by using \pagecolor again. Just the order (first \pagecolor before \usepackage{pdfpages}) is important.

3 Required Packages

The pdfpages package requires the following packages:

eso-pic CTAN:macros/latex/contrib/supported/eso-pic/
everyshi CTAN:macros/latex/contrib/supported/ms/

graphicx, ifthen, calc These packages belong to the standard L^{A}TEX distribution.

Furthermore it requires a recent version of:

pdftex.def http://www.tug.org/applications/pdftex/pdftex.def

Since pdfTEX, Version 3.14159-1.00a-pretest-20010806, PDF import has improved a lot. This results in much smaller file sizes, faster processing and the intuitively correct treatment of landscape pages. The latest version of pdfTEX can be found

at: ftp://ftp.muni.cz/pub/tex/local/cstug/thanh/pdftex.

4 Acknowledgment

I would like to thank Rolf Niepraschk and Heiko Oberdiek for their useful hints and suggestions. As well as Ross Moore, who encouraged me to implement the hypertext features.

References

[1] Hàn Thế Thành, Sebastian Rahtz, Hans Hagen, The pdfTEX user manual, http://www.tug.org/applications/pdftex/pdftex-s.pdf

[2] PDF Reference, Third Edition, Version 1.4, Adobe Systems Incorporated, http://partners.adobe.com/asn/developer/technotes/acrobatpdf.html